solve x^2=20, where x is a real number, simplify your answer as much as possible if mor than one solution separate
Answer by user_dude2008(1862)

x^2=20

x=sqrt(20) or x=-sqrt(20)


x=2*sqrt(5) or x=-2*sqrt(5)